OpenStudy (anonymous):

If y= -4/5x - 2, what is the value of x if y= 9?

OpenStudy (anonymous):

How would I solve this?

OpenStudy (anonymous):

Well you plug in y and rearrange the equation to isolate x.

OpenStudy (anonymous):

or plug in y and set the equation equal to zero

OpenStudy (anonymous):

0=(-4/5)x-2-9 0=(-4/5)x-11 11=(-4/5)x 13.75=x

OpenStudy (anonymous):

I think that should work.